# Going beyond density functional theory accuracy: Leveraging experimental data to refine pre-trained machine learning interatomic potentials

### v1

@@ Abstract
Machine learning interatomic potentials (MLIPs) are inherently limited by the accuracy of the training data, usually consisting of energies and forces obtained from quantum mechanical calculations, such as density functional theory (DFT). Since DFT itself is based on several approximations, MLIPs may inherit systematic errors that lead to discrepancies with experimental data. In this paper, we use a trajectory re-weighting technique to refine DFT pre-trained MLIPs to match the target experimental Extended X-ray Absorption Fine Structure (EXAFS) spectra. EXAFS spectra are sensitive to the local structural environment around an absorbing atom. Thus, refining an MLIP to improve agreement with experimental EXAFS spectra also improves the MLIP prediction of other structural properties that are not directly involved in the refinement process. We combine this re-weighting technique with transfer learning and a minimal number of training epochs to avoid overfitting to the limited experimental data. The refinement approach demonstrates significant improvement for two MLIPs reported in previous work, one for an established nuclear fuel: uranium dioxide (UO2) and second one for a nuclear fuel candidate: uranium mononitride (UN). We validate the effectiveness of our approach by comparing the results obtained from the original (unrefined) DFT-based MLIP and the EXAFS-refined MLIP across various properties, such as lattice parameters, bulk modulus, heat capacity, point defect energies, elastic constants, phonon dispersion spectra, and diffusion coefficients. An accurate MLIP for nuclear fuels is extremely beneficial as it enables reliable atomistic simulation, which greatly reduces the need for large number of expensive and inherently dangerous experimental nuclear integral tests, traditionally required for the qualification of efficient and resilient fuel candidates.
@@ Introduction
Molecular dynamics (MD) is an essential tool for many advancements in the fields of material science, biophysics, and chemistry.[1] However, the quality of MD simulations depends almost entirely on the accuracy of the interatomic potential used to calculate atomic forces. Although classical force fields offer the benefit of computational efficiency for predicting atomic forces, their accuracy is limited by the physics-based functional form and empirical model parameters that are fit to limited experimental data and/or quantum mechanical calculations[2]. In contrast, quantum mechanical calculations of atomic forces are accurate and transferable to a wide range of systems. However, these calculations are extremely computationally expensive, limiting their applicability to smaller systems.
@@
A possible solution to this fundamental limitation is to train directly to experimental data[14, 15, 16]. Classical and coarse-grained potentials are routinely parameterized with experimental data and, thus, tend to outperform MLIPs when tested against similar experimental observables. Leveraging experimental data to systematically construct MLIPs is not as straightforward as training to a quantum mechanical dataset[17]. Typically, an MLIP is trained using well established techniques, such as gradient descent, in which the atomistic positions are mapped to the corresponding target energies and forces. Such atomistic details are not measurable experimentally. Likewise, the experimental observables cannot be predicted directly from the MLIP. Rather, estimating experimental observables for an MLIP requires performing MD simulations and computing statistical averages of the instantaneous value for the macroscopic property over the MD trajectory. Furthermore, the measured quantities are susceptible to both experimental and simulation fluctuations, making training more challenging.
@@
Training to experimental data has previously been explored using automatic differentiation[18, 19, 20], in which, the model parameters are updated at the repetition of forward and backward passes. In the forward pass, an MD simulation is performed and observables are calculated. A loss function based on the difference between the predicted and target MD observables is defined. The gradients of the loss function with respect to the potential energy parameters are calculated in the backward pass. This approach suffers from three major problems: computational inefficiency, memory overflow, and explosive gradients[21].
@@
The differential trajectory re-weighting method aims to overcome the challenges of automatic differentiation[14]. In this approach, instead of calculating a new trajectory, a reference trajectory is re-weighted at every update of the MLIP parameters. Moreover, the functional relationship between observables and MLIP parameters established by re-weighting eliminates the calculation of back-propagation gradient at every step. This approach also allows for simultaneous training to both quantum mechanical data and experimental observables[15]. Previous work applying this technique demonstrates that the trained MLIP achieves good agreement with both the experimental observables included in the training as well as observables not included in the training dataset.[14, 15, 22, 23, 24].
@@
In this work, we use an approach similar to differential trajectory re-weighting where the statistical average for a set of simulation atomistic configurations are re-weighted in accordance with thermodynamic perturbation theory. However, in contrast to differential trajectory re-weighting, the proposed methodology does not require auto-differentiation of the macroscopic observable with respect to the MLIP parameters, rendering the retraining extremely fast and more straightforward. We refine the MLIP indirectly with the experimental spectroscopic observable known as Extended X-ray Absorption Fine Structure (EXAFS) spectra. EXAFS is an ideal property for training because of its high sensitivity to the local structural environment around an absorbing atom. However, our re-weighting technique is amenable to any thermophysical property that can be computed for a given atomic configuration, e.g., radial distribution functions. To avoid overfitting to a single experimental spectrum, the re-weighting technique is combined with transfer learning and re-training is performed for a limited number of epochs.
@@
We apply our refinement method to improve pre-trained MLIPs for an established nuclear fuel UO2 and a nuclear fuel candidate UN. Developing the interatomic potential models for nuclear fuels is particularly important because performing experiments is challenging due to their radioactive nature. Accurate simulations are crucial to complement the limited experimental data needed for testing fuel design, safety, and efficiency. Moreover, quantum mechanical calculations are extremely expensive for -electron elements, since the computational cost of traditional Kohn-Sham DFT method scales cubically with the system size,[25, 26] motivating the need for accurate MLIPs for these materials. Recently, MLIPs were developed for uranium dioxide (UO2)[27, 28] and uranium mononitride (UN)[29] using an active learning procedure to generate a DFT dataset. The UO2 MLIP was further improved by transfer learning to a smaller DFT+U dataset. The Hubbard parameter correction term (+U) improves the accuracy of DFT by accounting for the strong correlation among uranium 5 electrons. Both the DFT+U-based UO2 MLIP and the DFT-based UN MLIP provide a reasonable prediction of point defect energies, elastic constants, and other thermophysical properties over a range of temperatures. However, certain properties are poorly predicted by the respective MLIPs. For example, the UO2 MLIP quantitatively overestimates the lattice parameter compared to the experimental value. This overprediction in lattice parameter is a well-known limitation of the DFT+U method[30]. To a lesser extent, the UN MLIP underestimates the lattice parameter compared to the experimental value. However, the MLIP lattice parameter is in close agreement with the DFT value.[31] Thus, the deviations in lattice parameter with respect to experiment for both the UO2 and UN MLIPs can clearly be attributed to a deficiency in the DFT/DFT+U reference data, and not to a lack of diversity in the training dataset or the MLIP architecture choice.
@@
In this work, unrefined MLIPs for UO2 and UN are first trained to the corresponding reference DFT+U[27] or DFT datasets[29]. These MLIPs are then refined using our re-weighting technique and synthetic EXAFS data. The improvement in the refined MLIPs is validated by comparing the predicted thermophysical, crystal, energetic, and transport properties with experimental values. In particular, the refined MLIPs demonstrate excellent agreement with the experimental lattice parameter over a wide temperature range.
@@
The unrefined MLIP for UO2 was pre-trained on a DFT+U dataset consisting of 3365 structures,[27] while the unrefined MLIP for UN was pre-trained on a DFT dataset consisting of 12336 structures.[29] The respective training datasets for UO2 and UN were both generated previously using active learning (AL).[8] AL is an automated technique that ensures adequate coverage of the entire multidimensional potential energy surface while also reducing the size of the dataset by selecting only the most informative atomic configurations. For UO2, standard DFT calculations yield a metallic system in the ground state, in contrast to the insulator property known experimentally.[32, 33] Adding the Hubbard parameter (+U) to the exchange correlation functional in DFT calculations of UO2 correctly captures both the insulator behavior and the antiferromagnetic character. However, DFT+U vastly overpredicts the lattice parameters for UO2 compared to experiment. By contrast, standard DFT calculations are adequate and preferable for UN, as DFT+U yields negative phonon frequencies.[29]
Both MLIPs are trained with the Hierarchical Interacting Particle Neural Network (HIP-NN)[34], a message-passing graph neural network architecture (see FigureΒ 1). HIP-NN receives as inputs the element types and interatomic distances within the cutoff radius. These inputs are converted into a feature vector () of the local chemical environment around atom . The feature vector for the input layer is simply the one-hot encoding representation of atom with length equal to the number of element types, while the feature vector for all subsequent layers is of length for . The feature vector is passed through a series of interaction blocks, consisting of an interaction layer followed by atomic layers. In the interaction layer, the feature vector of the central atom is βmixedβ with the feature vectors for all of the neighboring atoms within a given cutoff radius.
@@
The HIP-NN hyperparameters for the UO2 MLIP are tuned to minimize the energy and force RMSE values, as detailed in Supporting Information. The selected hyperparameters for the UO2 MLIP are also listed in Supporting Information. A comparison is made between different loss functions by evaluating the bulk modulus for UO2 obtained using the resulting MLIPs (see Figure S1). In the case of UN, the same MLIP hyperparameters and loss function reported in previous work are used[29].
@@
Retraining an MLIP to experimental EXAFS spectra requires a spectrum for all elements in the system. Preliminary attempts to retrain our UO2 MLIP with an EXAFS spectrum for only U atoms resulted in significant degradation of the MLIP (see Figures S2 and S3). Unfortunately, experimental EXAFS spectra are only available for U atoms in both UO2 and UN. For this reason, we develop synthetic EXAFS spectra for each element, i.e., a separate spectrum for U and O in UO2, and for U and N in UN. These synthetic spectra serve as target macroscopic observables in the refinement procedure of their corresponding MLIPs. Each synthetic spectrum is generated for a perfect crystal modeled with the experimental lattice parameter (Figure S4). The experimental lattice parameter is used to ensure accurate interatomic distances and coordination shells for the perfect crystal and, consequently, to improve agreement between the synthetic EXAFS spectra and the experimental spectra. Generating synthetic spectra would not be necessary for systems where experimental EXAFS spectra are available for each element.
A comparison between synthetic and experimental EXAFS spectra for U in UO2 at 77 K is shown in Figure S5. The peak positions of the EXAFS spectra along the wavevector align well. Thus, the oscillations of the synthetic EXAFS spectra are consistent with the experimental spectra. However, the perfect crystal does not consider thermal vibrations or non-crystallinity present in the experimental measurements. Thermal effects dampen the spectral peaks, as the EXAFS amplitude is directly proportional to the Debye-Waller factor (second term in EquationΒ 4), which decreases exponentially with increasing mean square displacement of atoms .[40] For this reason, the peaks in the experimental EXAFS spectra are dampened compared to the perfect crystal (see Figure S5). To mitigate thermal effects, the MLIP refinement procedure is employed to match the synthetic EXAFS spectra at low temperatures- 75 K for UO2 and 50 K for UN. Preliminary attempts to refine the MLIP with EXAFS spectra at elevated temperatures were unsuccessful (see Figures S6 and S7).
@@
FigureΒ 1 shows the detailed MLIP refinement workflow. During this process, the MLIP pre-trained to the DFT/DFT+U dataset is iteratively re-trained to improve agreement with the target EXAFS spectra. Rather than re-training the unrefined MLIP directly to EXAFS data, our refinement methodology utilizes the EXAFS spectra to generate a new refinement dataset consisting of shifted atomic energies.
This refinement dataset is produced using the trajectory re-weighting technique in a multi-step process. The first step is to perform a set of MD simulations with the current MLIP. In the iteration, the current MLIP is the unrefined model pre-trained to DFT/DFT+U data. Snapshots, i.e., atomic configurations, from the MD trajectory are selected at equal time intervals. The atomic energies () for each snapshot are computed using the current MLIP. In the second step, EXAFS spectra are obtained for every atom in each snapshot. Third, these simulated EXAFS spectra are re-weighted to better match the target EXAFS spectra by minimizing a loss function (see EquationΒ 5). The component in the optimal re-weighting vector () is used to calculate atomic energy shifts , see EquationΒ 7 and EquationΒ 8. The refinement dataset consists of atomic configurations and shifted atomic energies for every atom in each snapshot.
Notably, the refinement dataset does not contain atomic forces. Training with forces typically results in more robust MLIPs than simply training to total energy,[41] for at least two key reasons. First, whereas total energy is a global property, atomic forces are a local property and, thus, are more rich in information content regarding the interactions governing a given atomic environment. Second, while each snapshot is assigned a single total energy, every atom contains a 3-dimensional vector of forces, greatly increasing the shear amount of data compared to a dataset consisting of solely total energies. Training to atomic energies is, thus, more similar to training to atomic forces because atomic energies are a local property and every atom is assigned an atomic energy.
@@
where is a matrix of simulation spectra and the re-weighting vector, , assigns each simulation spectrum a weight between zero and one. Inspired by thermodynamic perturbation theory,[42] the re-weighting vector is defined as:
@@
Thermodynamic perturbation theory, and by extension our re-weighting technique, is reliable only for small updates to the MLIP, or shifts to the atomic energy[43]. Large shifts in atomic energy lead to a strong imbalance in the re-weighting vector, where only a few configurations are assigned non-negligible weights. An imbalanced re-weighting vector results in poor statistical averaging and, thus, poor estimates of the re-weighted spectrum. Large shifts in atomic energy can also cause the refined MLIP to completely forget the qualitative landscape of the potential energy surface from the DFT/DFT+U dataset. For these reasons, a regularization term is added to the loss function (2nd term in Equation 5) to ensure that the absolute value of the atomic energy shifts are closer to zero. However, a balance in is necessary to ensure reliable estimates of the re-weighted spectra while also significantly improving the MLIP. To penalize large absolute values in while also allowing for meaningful shifts in the atomic energy, the regularization factor is chosen to be the smallest value that maintains the following inequality constraint:
@@
where for and for strongly imbalanced . The loss function (EquationΒ 5) is minimized using the optimization algorithm LM-BFGS (Limited-memory BroydenβFletcherβGoldfarbβShanno).[44]

### v2

@@ Abstract
Machine learning interatomic potentials (MLIPs) are inherently limited by the accuracy of the training data, usually consisting of energies and forces obtained from quantum mechanical calculations, such as density functional theory (DFT). Since DFT itself is based on several approximations, MLIPs may inherit systematic errors that lead to discrepancies with experimental data. In this paper, we use a trajectory re-weighting technique to refine DFT pre-trained MLIPs to match the target experimental Extended X-ray Absorption Fine Structure (EXAFS) spectra. EXAFS spectra are sensitive to the local structural environment around an absorbing atom. Thus, refining an MLIP to improve agreement with experimental EXAFS spectra also improves the MLIP prediction of other structural properties that are not directly involved in the refinement process. We combine this re-weighting technique with transfer learning to avoid overfitting to the limited experimental data. The refinement approach demonstrates significant improvement for two MLIPs, one for an established nuclear fuel: uranium dioxide (UO2) and second one for a nuclear fuel candidate: uranium mononitride (UN). We validate the effectiveness of our approach by comparing the results obtained from the original (unrefined) DFT-based MLIP and the EXAFS-refined MLIP across various properties, such as lattice parameters, bulk modulus, heat capacity, point defect energies, elastic constants, phonon dispersion spectra, and diffusion coefficients. An accurate MLIP for nuclear fuels is extremely beneficial as it enables reliable atomistic simulation, which greatly reduces the need for large number of expensive and inherently dangerous experimental nuclear integral tests, traditionally required for the qualification of efficient and resilient fuel candidates.
@@ Introduction
Molecular dynamics (MD) is an essential tool for many advancements in the fields of material science, biophysics, and chemistry [26]. However, the quality of MD simulations depends almost entirely on the accuracy of the interatomic potential used to calculate atomic forces. Although classical force fields offer the benefit of computational efficiency for predicting atomic forces, their accuracy is limited by the physics-based functional form and empirical model parameters that are fit to limited experimental data and/or quantum mechanical calculations [17]. In contrast, quantum mechanical calculations of atomic forces are accurate and transferable to a wide range of systems. However, these calculations are extremely computationally expensive, limiting their applicability to smaller systems.
@@
A possible solution to this fundamental limitation is to train directly to experimental data [60, Röcken2024, 39]. Classical and coarse-grained potentials are routinely parameterized with experimental data and, thus, tend to outperform MLIPs when tested against similar experimental observables. Leveraging experimental data to systematically construct MLIPs is not as straightforward as training to a quantum mechanical dataset [19]. Typically, an MLIP is trained using well established techniques, such as gradient descent, in which the atomistic positions are mapped to the corresponding target energies and forces. Such atomistic details are not measurable experimentally. Likewise, the experimental observables cannot be predicted directly from the MLIP. Rather, estimating experimental observables for an MLIP requires performing MD simulations and computing statistical averages of the instantaneous value for the macroscopic property over the MD trajectory. Furthermore, the measured quantities are susceptible to both experimental and simulation fluctuations, making training more challenging.
@@
Training to experimental data has previously been explored using automatic differentiation [4, 51, 66], in which the model parameters are updated at the repetition of forward and backward passes. In the forward pass, an MD simulation is performed and observables are calculated. A loss function based on the difference between the predicted and target MD observables is defined. The gradients of the loss function with respect to the potential energy parameters are calculated in the backward pass. This approach suffers from three major problems: computational inefficiency, memory overflow, and explosive gradients[8].
@@
The differential trajectory re-weighting method aims to overcome the challenges of automatic differentiation [60]. In this approach, instead of calculating a new trajectory, a reference trajectory is re-weighted at every update of the MLIP parameters. Moreover, the functional relationship between observables and MLIP parameters established by re-weighting eliminates the calculation of back-propagation gradient at every step. This approach also allows for simultaneous training to both quantum mechanical data and experimental observables [Röcken2024]. Previous work applying this technique demonstrates that the trained MLIP achieves good agreement with both the experimental observables included in the training as well as observables not included in the training dataset [60, Röcken2024, 35, 21, 18].
Another approach for refining an MLIP to experimental data is the density alignment method [20]. For this method, an MD simulation with constant temperature and pressure is performed with the pre-trained MLIP. Atomic positions and instantaneous box sizes are stored periodically. The deviation between the pre-trained MLIP simulation density for a given snapshot and the experimental density is included in the loss function. By assuming isothermal compressibility and applying the virial theorem for pressure in terms of the atomic forces, the derivative of density with respect to the MLIP parameters can then be derived with automatic differentiation. Thus, a key advantage of the density alignment method is the ability to apply standard backpropagation to train the MLIP parameters directly to experimental density data. The primary disadvantage of this method is that it is currently limited to density, although similar techniques may be applicable to other properties.
@@
In this work, we use an approach similar to differential trajectory re-weighting where the statistical average for a set of simulation atomistic configurations are re-weighted in accordance with thermodynamic perturbation theory. However, in contrast to differential trajectory re-weighting, the proposed methodology does not require auto-differentiation of the macroscopic observable with respect to the MLIP parameters, rendering the retraining extremely fast and more straightforward. We refine the MLIP indirectly with the experimental spectroscopic observable known as Extended X-ray Absorption Fine Structure (EXAFS) spectra. EXAFS is an ideal property for training because of its high sensitivity to the local structural environment around an absorbing atom. However, our re-weighting technique is amenable to any thermophysical property that can be computed for a given atomic configuration, e.g., the radial distribution function (RDF). To avoid overfitting to a single experimental spectrum, the re-weighting technique is combined with transfer learning.
We apply our refinement method to improve pre-trained MLIPs for an established nuclear fuel UO2 and a nuclear fuel candidate UN. Accurate modeling of these materials can significantly improve mechanistic and fuel performance models, reducing the number of challenging irradiation experiments needed to inform the design and licensing of new reactor systems and fuels. Developing the interatomic potential models for nuclear fuels is particularly important because performing experiments is challenging due to their radioactive nature. Moreover, quantum mechanical calculations are extremely expensive for -electron elements, since the computational cost of traditional Kohn-Sham DFT method scales cubically with the system size [32, 25], motivating the need for accurate MLIPs for these materials.
Recently, MLIPs were developed for uranium dioxide (UO2) [58, 12] and uranium mononitride (UN) [1] using an active learning procedure to generate the training datasets. As the accuracy of an MLIP is limited by the accuracy of the underlying reference data, these previous studies considered several different DFT functionals and values for the Hubbard parameter (+U), which is often added to the exchange correlation functional to account for the strong correlation among uranium 5 electrons. Because standard DFT calculations of UO2 yield a metallic system in the ground state, in contrast to the insulator property known experimentally [10, 13], the UO2 reference data were computed using DFT+U, which correctly captures both the insulator behavior and the antiferromagnetic character of UO2 [58, 12]. By contrast, the UN reference data were computed using standard DFT calculations, as DFT+U yields negative phonon frequencies for UN [1].
@@
Both the DFT+U-based UO2 MLIP of Stippell et al. [58] and the DFT-based UN MLIP of Alzate-Vargas et al. [1] provide a reasonable prediction of point defect energies, elastic constants, and other thermophysical properties over a range of temperatures. However, certain properties are poorly predicted by the respective MLIPs. For example, the UO2 MLIP quantitatively overestimates the lattice parameter by compared to the experimental value. This overprediction in lattice parameter is a well-known limitation of the DFT+U method[7]. To a lesser extent, the UN MLIP underestimates the lattice parameter by compared to the experimental value. However, the MLIP lattice parameter is in close agreement with the DFT value[31]. Thus, the deviations in lattice parameter with respect to experiment for both the UO2 and UN MLIPs can clearly be attributed to a deficiency in the DFT+U/DFT reference data, and not to a lack of diversity in the training dataset or the MLIP architecture.
@@
In this work, new unrefined MLIPs for UO2 and UN are first trained to the corresponding reference DFT+U[58] or DFT datasets[1]. These MLIPs are then refined using our re-weighting technique and synthetic EXAFS data. The improvement in the refined MLIPs is validated by comparing the predicted thermophysical, crystal, energetic, and transport properties with experimental values. In particular, the refined MLIPs demonstrate excellent agreement with the experimental lattice parameter over a wide temperature range, clearly achieving beyond DFT-level accuracy.
@@
The unrefined MLIP for UO2 was pre-trained on a DFT+U dataset consisting of 3365 structures[58], while the unrefined MLIP for UN was pre-trained on a DFT dataset consisting of 12336 structures[1]. See the literature for details regarding the DFT+U/DFT calculations (e.g., functional, Hubbard parameter).
Both MLIPs are trained with the Hierarchical Interacting Particle Neural Network (HIP-NN) [37], a message-passing graph neural network architecture (see Figure 1). HIP-NN receives as inputs the element types and interatomic distances within the cutoff radius. These inputs are converted into a feature vector () of the local chemical environment around atom . The feature vector for the input layer is simply the one-hot encoding representation of atom with length equal to the number of element types, while the feature vector for all subsequent layers is of length for . The feature vector is passed through a series of interaction blocks, consisting of an interaction layer followed by atomic layers. In the interaction layer, the feature vector of the central atom is “mixed” with the feature vectors for all of the neighboring atoms within a given cutoff radius.
@@
The HIP-NN hyperparameters for the UO2 MLIP are tuned to minimize the energy and force RMSE values (see Tables S1, S2, and S3). A comparison is made between different loss functions by evaluating the bulk modulus for UO2 obtained using the resulting MLIPs (see Table S4 and Figure S1). The selected hyperparameters for the UO2 MLIP are included in the Supporting Information. For UN, the same MLIP hyperparameters and loss function reported in previous work are used[1].
@@
Retraining an MLIP to experimental EXAFS spectra requires a spectrum for all elements in the system. Preliminary attempts to retrain our UO2 MLIP with an EXAFS spectrum for only U atoms (see Figure S2) resulted in significant degradation of the MLIP (see Figure S3 and the next section for a detailed explanation). Unfortunately, for both UO2 and UN, experimental EXAFS spectra are only available for U atoms. Therefore, we develop synthetic EXAFS spectra for each element, i.e., a separate spectrum for U and O in UO2, and for U and N in UN. These synthetic spectra serve as target macroscopic observables in the refinement procedure of their corresponding MLIPs. Each synthetic spectrum is generated for a perfect crystal modeled with the experimental lattice parameter (see Figure S4). The experimental lattice parameter is used to ensure accurate interatomic distances and coordination shells for the perfect crystal and, consequently, to improve agreement between the synthetic EXAFS spectra and the experimental spectra. Generating synthetic spectra would not be necessary for systems where experimental EXAFS spectra are available for all elements. EXAFS spectra are needed for all elements even when using synthetic spectra, as evidenced by the sporadic behavior of the refined MLIP at high temperatures if the refinement process is performed with synthetic spectra for either U or O (see Figure S5).
A comparison between uranium synthetic and experimental EXAFS spectra in UO2 at 77 K is shown in Figure S6. The peak positions of the EXAFS spectra along the wavevector align well. Thus, the oscillations of the synthetic EXAFS spectra are consistent with the experimental spectra. However, the perfect crystal does not consider thermal vibrations or non-crystallinity present in the experimental measurements. Thermal effects dampen the spectral peaks, as the EXAFS amplitude is directly proportional to the Debye-Waller factor (second term in Equation 4), which decreases exponentially with increasing mean square displacement of atoms [65]. For this reason, the peaks in the experimental EXAFS spectra are dampened compared to the perfect crystal (see Figure S6). To mitigate thermal effects, the MLIP refinement procedure is employed to match the synthetic EXAFS spectra at low temperatures 75 K for UO2 and 50 K for UN. Preliminary attempts to refine the MLIP with experimental EXAFS spectra for U at higher temperatures (300 K) resulted in an unstable MLIP (see Figures S7 and S8).
In general, the refinement process does not necessarily require synthetic target spectra, as long as experimental data are available for all elements. We provide a simple example where experimental spectra are available for pure aluminum (see Appendix). In this example, the pre-trained MLIP is refined using RDF data at two temperatures, demonstrating that the refinement process is not limited to EXAFS data and is applicable to multiple temperatures.
@@
Figure 1 shows the detailed MLIP refinement workflow. During this process, the MLIP pre-trained to the DFT/DFT+U dataset is iteratively re-trained to improve agreement with the target EXAFS spectra. Rather than re-training the unrefined MLIP directly to EXAFS data, our refinement methodology utilizes the EXAFS spectra to generate a new refinement dataset consisting of shifted atomic energies.
The refinement dataset is produced using the trajectory re-weighting technique in a multi-step process. The first step is to perform a set of MD simulations with the current MLIP. In the iteration, the current MLIP is the unrefined model pre-trained to DFT/DFT+U data. Snapshots, i.e., atomic configurations, from the MD trajectory are selected at equal time intervals. The atomic energies () for each snapshot are computed using the current MLIP. In the second step, EXAFS spectra are obtained for every atom in each snapshot. Third, these simulated EXAFS spectra are re-weighted to better match the target EXAFS spectra by minimizing a loss function (see Equation 5). The component in the optimal re-weighting vector () is used to calculate atomic energy shifts , see Equation 7 and Equation 8. The refinement dataset consists of atomic configurations and shifted atomic energies for every atom in each snapshot. Because is computed only for elements with target EXAFS spectra, the atomic energies for elements without target EXAFS spectra would remain unshifted (), resulting in a data inconsistency when EXAFS spectra are not available for all elements.
Notably,because the original DFT/DFT+U forces are not compatible with the shifted atomic energies, the refinement dataset does not contain atomic forces. There are at least two key reasons why training with forces typically results in more robust MLIPs than simply training to total energy [54]. First, whereas total energy is a global property, atomic forces are a local property and, thus, are more rich in information content regarding the interactions governing a given atomic environment. Second, while each snapshot is assigned a single total energy, every atom contains a 3-dimensional vector of forces, greatly increasing the shear amount of data compared to a dataset consisting of solely total energies. Training to atomic energies is, thus, more similar to training to atomic forces because atomic energies are a local property and every atom is assigned an atomic energy.
@@
where is a matrix of simulation spectra and the re-weighting vector, , assigns each simulation spectrum a weight between zero and one. Inspired by thermodynamic perturbation theory [71], the re-weighting vector is defined as:
@@
Thermodynamic perturbation theory, and by extension our re-weighting technique, is reliable only for small updates to the MLIP, or shifts to the atomic energy [41, 42]. Large shifts in atomic energy lead to a strong imbalance in the re-weighting vector, where only a few configurations are assigned non-negligible weights. An imbalanced re-weighting vector results in poor statistical averaging and, thus, poor estimates of the re-weighted spectrum. Large shifts in atomic energy can also cause the refined MLIP to completely forget the qualitative landscape of the potential energy surface from the DFT/DFT+U dataset. For these reasons, a regularization term is added to the loss function (2nd term in Equation 5) to ensure that the absolute value of the atomic energy shifts are closer to zero. However, a balance in is necessary to ensure reliable estimates of the re-weighted spectra while also significantly improving the MLIP. To penalize large absolute values in while also allowing for meaningful shifts in the atomic energy, the regularization factor is chosen to be the smallest value that maintains the following inequality constraint:
@@
where for and for strongly imbalanced . The loss function (Equation 5) is minimized using the optimization algorithm LM-BFGS (Limited-memory Broyden–Fletcher–Goldfarb–Shanno) [36].